OpenStudy (anonymous):

Which of the following triangles is the right triangle? Answer A. 4, 4, 4 B. 18, 19, 35 C. 16, 63, 65 D. 9, 12, 20

OpenStudy (anonymous):

If \(c^{2} = a^{2} + b^2\), then it's a right triangle. If \(c^2 > a^2 + b^2\), then it's an obtuse triangle. If \(c^2 < a^2 + b^2\), then it's an acute triangle.

OpenStudy (anonymous):

Use the Pythagorean Theorem which is a^2 +b^2=c^2 c=the largest number in the set. a + b are the other two numbers. *if the numbers equal each other on both sides of the equation it is the right triangle.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

np :)